14.A company has 5-year bonds outstanding that pay an 7.5 percent coupon rate. Investors buying the bond today can expect to earn a yield to maturity of 14.4 percent p.a.. What should the company's bonds be priced at today? Assume annual coupon payments and a face value of $1000. (Rounded to the nearest dollar)

Select one:

a. $765

b. $1279

c. $638

d. $1959

15.Jack is planning to buy a 9-year bond with semi-annual coupons and a coupon rate of 4.3 percent p.a. The face value is $1,000. Given an annual yield of 6.4 percent, what is the bond’s current price? (to the nearest cent)

Select one:

a. $858.00

b. $1154.03

c. $572.17

d. $859.62

16.Which ONE of the following statements is NOT true?

Select one:

A. Long-term bonds have lower price volatility than short-term bonds.

B. As interest rates decline, the prices of bonds increase; and as interest rates rise, the prices of bonds decline.

C. All other things being equal, short-term bonds are less risky than long-term bonds.

D. Interest rate risk increases as maturity increases.

17.Jill wants to buy 5-year zero coupon bonds with a face value of $1,000. Her required return on the bonds is 6.9 percent p.a. Assuming annual compounding, what would Jill be prepared to pay for the bond? ( to the nearest cent )

Select one:

a. $979.44

b. $1020.83

c. $716.33

d. $847.89
Answer #1

Answer 14:

Correct answer is:

a. $765

Explanation:

Par value = $1,000

Annual coupon amount = 1000 * 7.5% = $75

Time to maturity = 5 years

Price of bond now = PV (rate, nper, pmt, fv, type) = PV (14.4%, 5, -75, -1000, 0) = $765

Option a is correct and other options b, c and d are incorrect.

Answer 15:

Correct answer is:

a. $858.00

Explanation:

Par value = $1,000

Semiannual coupon amount = 1000 * 4.3% = $21.50

Time to maturity in semiannual periods = 9 * 2 = 18

Price of bond now = PV (rate, nper, pmt, fv, type) = PV (6.4%/2, 18, -21.50, -1000, 0) = $858.00

Option a is correct and other options b, c and d are incorrect.

Answer 16:

Correct answer is:

A. Long-term bonds have lower price volatility than short-term bonds.

Explanation:

Price volatility is equal to yield times its duration. So higher the duration higher the price volatility. The statement A is incorrect.

The other 3 statements are correct.

As interest rates decline, the prices of bonds increase; and as interest rates rise, the prices of bonds decline.

All other things being equal, short-term bonds are less risky than long-term bonds.

Interest rate risk increases as maturity increases.

Option a is correct and other options b, c and d are incorrect.

Answer 17:

Correct answer is:

c. $716.33

Explanation:

PV = FV / (1 + Periodic interest rate) ^Number of periods

Price Jill be prepared to pay for the bond = PV = 1000 / (1 + 6.9%) ^5 = $716.33

Option c is correct and other options a, b and d are incorrect.